On Aeternal Philosophy and Religious Beliefs

Eirien Lisándrie

September 2004

Introduction

Aeternal philosophy is a complicated subject, and rather different from traditional belief systems, such as Christianity and Buddhism. Not only can it be fundamentally different, but due to the uncountable number of Aeternals in existence, it can be easily proven with an number of counter-examples that the philosophy of the Aeternalae is hardly homogenous, and likewise with their religious beliefs. The intent of this document however, is to illustrate one of the more common philosophical and belief systems held by the Aeternalae, in particular that of the True Moon Conpurifa Army.

Similarities to Western Philosophy

The Aeternal religious beliefs are fundamentally similar to the now unfamiliar Gnostic Christian and Discordian philosophies. Indeed, the idea of gnosis, as it applies to the Aeternalae is often considered to be a return to the Third Realm, which is as unknowable as the traditional Pauline Judeo-Christian god is supposed to be. In fact, the Christian Jesus, commonly thought of as The Eternal King, can be viewed as a partial attempt in illustrating the true nature of the universe and man from an unknown source.

Discordian influence is considerably less than that of Gnosticism, though it is still present, in part thanks to the influence of Eris in the birth of the Aeternalae. Few Aeternals actually properly further the purposes of chaos in the universe directly (through such things as the Turkey Curse), though all Aeternals who actually bother to rescend First Realm worlds are indirectly supporting Erisian attempts to promulgate chaos as well as fighting the Gnostic Demiurge.

The Nature of the Universe

Aeternal philosophy fundamentally believes in three primary realms of the universe:

1. The Second Realm, which was the first realm to have ever existed in the universe. The method of creation of the Second Realm is debateable, and relatively unknown. The Second Realm now exists, however, as the reservoir of dreams, imagination, and energy for the entire universe.

2. The Third Realm, which was the second realm to exist, and the first of the Second Realm’s children. The Third Realm is home to many ethereal creatures, called demons, as well as the transient home of souls who have died in First Realm worlds.

3. The First Realms, which were the last realms to be created, and the last of the Second Realm’s children. The First Realms are home to what is commonly considered to be “reality” by non-demon and non-fae creatures, who inhabit it.

In simplified terms, the purpose of these realms can be stated as “matter, energy, and infinity” or “reality, possibility, and eternity” for the First, Second, and Third Realms respectively. In addition to these however, are also many smaller “subrealms” such as Rekai or Makai, which serve as the “glue” between the First and Third Realms directly, as well as safe-havens from the overbearing chaotic nature of the Third Realm on its own.

In addition, while the more basic glue by which travel between realms can be performed is not completely understood by human scientists, we are lucky enough to have discovered the general method: quantum mechanics. In all actuality, quantum mechanics, and the probability that it works on is actually only the visualization of such in the First Realm, as the idea of quantum mechanics as humanity knows it is actually inapplicable in the other two realms, due to physics as we know it not appearing in the same nature (though they still play a major role).

The First Realm

The First Realm, as previously described, is home to all “mortal” creatures such as humans and other ningen, and other less intelligent creatures, such as animals, plants, and other “living creatures” (as defined by human knowledge). In short, the First Realm is reality as First Realm inhabitants are limited to knowing, as the inner workings of the Second and Third Realms are, for all practical purposes, unknowable to human consciousness.

The First Realm is actually an infinite number of “real” universes, of which actually a very small number of worlds (though still infinite in number) actually contain living creatures. New First Realms are constantly being formed from the Second Realm, in part to offset the rescension and destruction of these worlds by the Aeternalae and other demonic creatures alike, but also in response to the virtual creation of imaginary worlds in the imaginations of other First Realm inhabitants. As a result, the “fictional” worlds of Middle-Earth, Riven, Discworld, the Milky Way of The Hitchhiker’s Guide to the Galaxy, and many other worlds actually do exist, though they are beyond reach of conventional First Realm technology. This is because First Realms are still completely selfsufficient universes that are impossible to jump between through normal means.

Furthermore, there are a large number of First Realms created as “alternate worlds” where they are merely worlds that branch off at a nexus point with one small change. As a result, all “alternate histories” that both have and have not been proposed by authors exist, with some being extremely similar to our current universe, having no significant changes in the long run. This is often what an Aeternal is referring to when stating that time is an illusion.

The Second Realm

The Second Realm is the realm of dreams, imagination, and energy. There are relatively few native inhabitants of the Second Realm, though they do exist, known commonly as the fae, the muses, and other similar beings. All those who imagine and create use the energy of the Second Realm to accomplish these goals. In addition, as the realm of dreams, all dreams are realized in it, and thus, while unknowing of it, the Second Realm is the only other realm that First Realm beings enter, if fleetingly. There are a few exceptions to this rule, being the more spiritually inclined tribal peoples who have learned long ago how to dreamwalk properly...

Third Realm beings, such as the Aeternalae are particularly known for dreamwalking and even using the energy of the Second Realm by way of dreamwalking to accomplish their goals.

The Third Realm

The Third Realm is also unknown to ningen thought, as it is the realm of demons, both good and bad. Unlike demons of Judeo-Christian culture, not all of these demons are necessarily evil, though many can be considered evil by those not exactly eager to receive their help.

The Third Realm has not physical existence, and it is infinite in boundaries, always just sparse enough for third realm beings to easily get around and have privacy, even when limited to “habitable” regions. In the Third Realm properly however, each being essentially lives in an isolated “universe” created as they wish, and interacting with others’ if they so desire. Time, space, and all kinds of dimensions lack meaning in the Third Realm, and thus in many ways the Third Realm represents a “paradise” or “hell” depending on who you ask.

The majority of the Third Realm, in fact, is pure chaos and undescribable, and thus actually rather unsuitable for beings to live in. As a result of the influence of such subrealms as Rekai and Makai however, the Third Realm can gain less chaotic aspects at certain points, thus allowing for the situations described above.

Subrealms

Rekai and Makai serve, as mentioned before, as primary “gateways” to First Realms from the Third Realm. The nature of these subrealms is not as well known as other realms, but they are similar, in many ways, to a First Realm in the Third Realm, that is, it has a consistent structure and social hierarchy. For as yet unknown reasons, the nature of the subrealms creates islands of stability inside the otherwise chaotic Third Realm, thus allowing for spiritual beings to leave the subrealms to live peacefully.

The division between Rekai and Makai is rather stark, in that it seems to follow the rather similar divisions of ningen concepts of “good” and “evil” despite the relativity of these statements. Because of this, some spiritual beings are refused access to either one or the other subrealm. Aeternalae are an exception to this rule, being given access to enter both Rekai and Makai. This closely parallels their confusing appearance to ningen as alternately “good” and “evil”. While there may be other subrealms like that of Rekai and Makai, their qualities are not as well known to the author.

Aeternal Philosophy

In order to better understand the nature of Aeternal religion, one must begin with a study of Aeternal philosophy. Aeternal philosophy is actually quite similar to post-modernist philosophy of recent years, promoting a more relative universe in which one must learn to adapt to life appropriately, considering life to be a game.

Absolute Truth

One of the primary Aeternal beliefs is thus the lack of an absolute truth, in contrast with that of Christianity. Truth and falsehood is all relative to the person who observes something that happens, not only from location in space and time, but also in perspective of personality and personal history, and as a result, all truth is inevitably tainted with a small bit of falsehood.

In a similar vein, Aeternals also often maintain that there is no absolute sense of right and wrong. While this may make it seem impossible to maintain justice in Aeternal affairs, justice is no longer an absolute, but rather it is defined by a group of people who share similar senses of right and wrong (whether it be an Aeternal faction, a culture, or something else). This can be carried to conflict, such as in the following common example:

Inevitably, any world that an Aeternal seriously tries to rescend will begin to resist efforts to be rescended for numerous reasons. For the most part, it is believed that the originae is forcing an undesireable goal on the people of the world being rescended, and thus conservative elements band together to fight against the impending onslaught. These conservative forces thus have a similar sense of justice that defines what an Aeternal is doing is “wrong” and that life in the First Realm is “right.” In contrast, the originae likely believes that it is right and just to bring First Realms into the fold of “truth” in the Third Realm (which we will examine in greater detail later). As a result of this, there grows a conflict between two senses of justice, boiling into a war between both sides that believe (and ARE right). A similar ideological conflict exists on Earth in the “war” between America and Islamic extremists. Aeternal philosophy maintains that all conflict can find its root in such inconsistencies between groups’ ideals of justice.

Destiny

If there is one thing that the Aeternalae DON’T believe, it is in the instrument of destiny. Indeed, destiny and karma (the latter at the least in a long-term sense) are generally conceived to be ningen interpreations of the mechanations of the Second Realm.

As Aeternalae regularly flaunt the rules of the Second Realm, it is understandable that destiny is regularly altered by them. In fact, the influence created after the creation of several new etones completely rewrites “destiny” as might be conceived in ningen thought. So while destiny exists, it is not as set in stone as ningen are regularly led to believe. Free will, though illusory on a certain level (after all, one is conditioned to react to certain things in certain ways) does exist on the larger scale.

As for karma however, the significance of such is diminished, due to the assumption of the universe that people are inherently good (rather than inherently evil or a blank slate, tabula rasa) but rather easily hurt and confused. As a result, people who are “evil” are not punished by some overriding cosmic force. Instead, they are to be healed, which is one thing that the Aeternalae strive to do, even if it is by their own methods.

Short-term karma, in particular, karma in this life, is not to be completely ruled out however. After all, it is a known fact that people are normally influenced by your own actions, and as such, the Golden Rule very much applies. so while good things often happen to good people, and bad things happen to bad people, it’s not necessarily due to predestination of their actions.

The Feminine Essence

If there is one belief that is most common among the Aeternalae, it is that almost all beings have the feminine spirit at their core, even if it is masked by a masculine appearance and action. Indeed, this belief is only strengthened by their ability to transform many ningen into their feminine ideal as exemplified in their Aeternal form. The Aeternals do acknowledge some exceptions to the rule, including beings who simply do not have the feminine ideal in them at all, as well as those beings who may have the feminine ideal, but simply have their souls designed in such a way as to prevent the overbearing destructive nature of Aeternalae rescencion due simply to the selfish desire to rescend worlds. Aeternals also belief that the feminine spirit is the most fundamental and important aspect of a person, and thus they are often moved to try to discard all layers that disguise a person from their true feminine essence: in other words, transforming people into Aeternals, or at the very least, female.

The Void

Like many Asian philosophies, Aeternals believe very much in the power of what they call the Void. The Void serves as a force of emptiness and non-existence. In contrast to Western dogma, non-existence is not viewed by the Aeternalae as necessarily an evil. Indeed, Buddhism serves very much as an example of positive Void energy in Earth culture. An example of the Void in practice is through meditation, as one creates emptiness inside the mind in order to more fully experience the complete nature of the universe as it exists.

The more destructive aspect of the Void is also called the Blood Void. The Blood Void is primarily concerned with complete and utter destruction to generate emptiness. The ultimate solution to suffering, according to the Zerouboros and others who follow the path of Blood Void, is by removing both that which causes suffering as well as that which suffers, thus hoping to achieve a negatively stable end to the universe by removing all that exists and returning the universe to nothingness.

Suffering

Another similarity to Asian philosophies and religions survives in the focus on the ending of suffering. While some Aeternals (such as the Yalapil and The Glorious Three Sun Army) consider ending the suffering of planets necessary to the point that they will cause immense temporary suffering to end a planet’s suffering, other groups consider that paths with minimal amounts of suffering are much preferred.

Thus, the matter of suffering and it’s definition are very important to Aeternals. The general definition is somewhat different and more comprehensive than the definition offered by Asian religions. While Asian religions and Aeternal philosophy agree that suffering can be caused by a difference between reality and desired fantasy, Aeternalae go one step further (taking from their Discordian roots) and also believe that suffering also can exist in forced stasis, especially when applied on a large scale, such as the attempts of ningen to try to enforce stasis upon Gaia such that they may continue to exist with their own selfish desires, causing even more suffering on the lower classes who also receive a large portion of this enforced stasis (see Orwell’s 1984).

The True Moon Conpurifa Army, recently exiled, is currently in the process of reforming its opinion of how best to offset suffering, and as such, it is regretful that the author cannot better elucidate the reader on Aeternal methods of ending suffering.

Aeternal Religion

Aeternal religions as a whole are completely disparate, as there are equal portions of Aeternals claiming pacificism as the correct way of life as those claiming for a need to conquer. Thus it can be said that there is no One True Aeternal Religion. Despite these apparently contradicting belief systems, there are many aspects of most Aeternal belief systems that are held in common.

Deities

One of the most prevalent aspects, both in their mythology and in reality, is the intimate relationship of deities with Aeternals. From the ancient mythos from which the Aeternals draw their heritage to the present, Aeternals and their methods, purposes, and plans have been fundamentally shaped by the actions of deities. Those held in particular reverence by the True Moon Conpurifa Army include Eris and Artemis, though other Alphaoriginae hold other deities unfamiliar to Earthlings as sacred.

Most deities were created long ago, as the first beings in the universe. Originally concepts given powers, deities have evolved in many different fashions. Several primary aspects of deities are common however, namely that their will is bound, they lack a proper soul present in most other beings, and they lack direct influence on the First and Third Realms. This is not to say that they are completely useless however, as their indirect influence, often through the guise of priests, priestesses, and fallible mortal “avatars,” is quite powerful.

Deities, however, lack free will, instead reacting as base, strictly logical beings. Their emotions are generally artificial, and their reactions are “preprogrammed” in a sense, being predictable. In this sense, deities can be possibly considered artificial beings, thus explaining the plethora of flat caricatures and archetypes present in Earth mythologies. In fact, the closest analogy for a deity would be that of an artificial intelligence program.

Because of this weakness in deities however, Aeternals tend to believe that deities are hardly anything to be feared or respected, as they are incapable of conscious decision-making. While they may inspire or offer advice, it is, at best, one-sided and predictable.

Eris, on the other hand, is an odd being, both above, below, and identical to a deity, and perhaps the only one of her kind. Eris is generally conceived of as being random, and having more direct dealings with the First and Third Realms, though still limited. Eris overall defies definition and cannot be clearly defined as either with or without will, being both inside and out of the system at the same time. The only thing that can be known for certain is that she opposes the force of the Demiurge (or Greyface, as he is called in the Principia Discordia).

Gnostic Elements in Aeternal Religion

With the Aeternalae’s purpose so strictly defined by their powers, it is unsurprising that their religions often have Gnostic elements supporting their ability to “rescend” worlds, even going so far as to justify it.

Gnosticism traditionally holds that the Monad, or the One Unknowable God, eminated many pairs of lesser Aeons who were also aspects of God. According to Gnostics, these Aeons together with the Monad make up the Pleroma or fullness of God. Sophia, paired with Christ, was a member of the lowest pair of the Aeons, and sought the Monad, according to Gnostic creation stories, and inadvertently created the world. In addition, in an attempt to eminate without Christ, she created the Demiurge, a lesser “God” focused only on the material objects and sensual pleasures of the world. As a result, Christ was forced to introduce both Sophia and mankind to spirit (the latter in the guise of Jesus) in order to enable mankind to return to the true God, the Monad through gnosis.

In the Aeternal version of this Gnostic legend fits the entire birth of the three realms, as well as the need to rescend. The Second Realm, by being first, and creating second the Third Realm, acts thus like the emination of Aeons to Aeternals, and the creation of First Realms is viewed analogous to the faulty creation of the world. Thus it is unsurprising that by rescending First Realms and introducing their inhabitants to their supposed female essence inside, the Aeternalae picture themselves forcing a hand in their idea of gnosis, by returning the physical world to the Second Realm, and introducing the inhabitants to the Third Realm.

While the Aeternalae have many natural enemies, simply based on their low stature as demons, and their attempts to further the known and existing universe, their greatest enemy is present in the spirit of the Demiurge. Unlike in classical Gnosticism, the Demiurge is not viewed as an actual being, but rather a spirit that often fills both demons and First Realm dwellers with love for the material world. Naturally, this impulse to love the material world and create to disastrous excess, as can be seen in some of Earth’s more egregious excesses, such as pollution, tends to work against Aeternals, as it leads to conservatism, and a desire to not leave the First Realm for the Third. Nevertheless, Aeternals often are seen as choosing to rescend worlds held strongly by the spirit of the Demiurge, in part to fight it, and also to enjoy their rescensions.

While not all Aeternals completely spurn a love for the material, all end up spurning at least the First Realm, relegating it as lost and unnecessary, other than for introducing others to their form. Indeed, by clinging to existence and the universe as it is, those of Zerouboros and others who support the Blood Void tend to claim hypocrisy in the Aeternal philosophy of rescension as they would prefer the universe to not exist at all, the extreme “logical” conclusion of Gnostic philosophy.

Aeternal Dogma

Among the rules common to Aeternalae are many aspects of Aeternal dogma. Indeed, the most widespread of these dogmatic principles and the one that sets them most at odd with Zerouboros is that of “love yourself first.”

Love Yourself First

The concept of loving one’s self first often strikes ningen as odd, arrogant, and selfish, and in many ways, it is. Despite this however, the concept of self-love is not all that arrogant. It is generally believed that loving one’s self tends to make one MORE compassionate towards others, rather than less. In context of a situation in which humanity would best understand, suppose the cabin of an aircraft depressurizes while in flight as a child is sitting next to you. While normal compassion would contrain one to help the child first, if one does not, and instead falls unconscious while doing so, one is actually less helpful, losing two lives instead of at most one. By helping one’s self first, one can thus be more equipped to help others by being in a better position to place the mask on the child. In this fashion, by considering one’s self first and foremost, a group benefits, rather than perhaps none.

In summary, Aeternals believe that those who cannot help themselves end up not helping anyone, and for this reason, one should love one’s self and help it first and foremost in order to better help others.

Live for the Moment

As a side-effect of the Aeternal concept of time is the fact that, as time is relative and thus not absolute in the least, the only time that can be well understood and thus be revered is the present, as the past is done and the future is malleable.

Because of this fact, the Aeternalae tend to accept a temporal belief based on “carpe diem” in many ways. If one cannot pin oneself in time, but can only say with conviction that time is moving forward, then the focus is naturally on the present, as one would not be very worried about what is ahead or behind.

This principle too can be extended even to ningen thought, particularly with regards to mortality, just as much as it can to immortality, though for alternate reasons. Thus, it is very important to release one’s self from worrying about what is already done and what will happen, as this necessarily distracts the mind from focusing on causality, which, to Aeternalae, is much more important than mere potentiality.

In addition, Aeternalae believe that the future does not come directly to someone, but in contrast, one moves towards the future, or alternately not at all. This proves to be an excellent example of the planetary stasis that is often imposed by and on populations that causes suffering. This naturally means that Aeternalae are very focused on pushing onwards to the future, rather than focusing so exclusively on the now to the point that they would not leave a point even when it becomes part of the past, as that naturally represents a sort of metaphysical attachment that is so representative of the Demiurge.

Conclusion

Ironically, while a Gnostic belief tends to justify rescension of First Realm worlds, most Aeternals tend to like the First Realm anyway, as it validates the existence of Aeternals and offers great challenges to those seeking to rescend. To other Aeternals, it also serves as a contact with the creative Second Realm through methods other than dreamwalking, allowing them to actually create in a sense otherwise impossible elsewhere. In a sense, it allows the Aeternalae to exist in context, and offering a reason to act and spread their feminine essence. Indeed, many Aeternals often pass through First Realms acting as noninterfering observers.